OpenStudy (anonymous):

-4g - 1/5= 4/5? with steps ? plz help i'm a dumb cookie :/

OpenStudy (anonymous):

1º add 1/5 bouth sides: -4g=1 2º devide by -4 bouth sides: g=-1/4

OpenStudy (anonymous):

i'm really confused ...

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

1º -4g -1/5 +1/5 =4/5 +1/5 -4g =5/5=1 2º -4g/-4 =1/-4 g=-1/4

OpenStudy (anonymous):

ohh :D

OpenStudy (anonymous):

thanks !!

OpenStudy (ganpat):

isolate g... use following steps...  can u solve for g ???